OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin2x + sin x = 0

sam (.sam.):

Do you know what's the identity for sin2x?

OpenStudy (anonymous):

noo i forgot..

sam (.sam.):

sin2x=2sinxcosx

sam (.sam.):

So how do you proceed?

OpenStudy (anonymous):

2sinx+cosx+sinx=0 right? what do i do next?

sam (.sam.):

no, it should be from sin2x + sin x = 0 to 2sinxcosx + sin x =0 right?

OpenStudy (anonymous):

ohh yea

sam (.sam.):

you should try factoring now

OpenStudy (anonymous):

sinx(2cosx+1)=0 is this right? what do i do next? o.O

OpenStudy (aravindg):

ya u r right :)

sam (.sam.):

yes, now you can let sinx=0 and 2cosx+1=0 then find the x values

OpenStudy (aravindg):

equate that to 0 and find x values

OpenStudy (anonymous):

so i got cosx=-1/2 how do i find the value of x?

sam (.sam.):

well you can do this cosx=-1/2 x=cos^{-1} (-1/2)

OpenStudy (anonymous):

so what are all possible values of x? 0 and......what else O.o

OpenStudy (aravindg):

@hannahg :) instead of o.O

sam (.sam.):

no its not 0

OpenStudy (anonymous):

@AravindG :)

sam (.sam.):

you calculate cos^(-1) 1/2 first without minus on half you'll get 60 degrees

OpenStudy (anonymous):

yes:)

sam (.sam.):

then use this chart for cosine  we are looking at the minus side since we got minus half. then 180-60=120 degrees 180+60=240 degrees

OpenStudy (aravindg):

@hannahg thats better :)

OpenStudy (anonymous):

yes. so all the possible solutions will be pi, pi/3, 2pi/3?

sam (.sam.):

you still have sinx=0 to do

OpenStudy (anonymous):

yes......! so 0?

sam (.sam.):

yes but dont forget about 180

sam (.sam.):



sam (.sam.):

since theta is 0, then you get 180

OpenStudy (anonymous):

@.Sam. so what are all solutions in the interval??